In children's park, there was a slide to be made by contract. By mistake, the person who had taken the contract made the coefficient of friction of the slide as high as 1/3. Now, the fun is that the child expecting to slide down the incline will stop somewhere in between. Find the angle θ with the horizontal at which he will stop he stopon the incline. (Assume negligible frictional losses)
1
Expert's answer
2020-07-30T10:38:55-0400
"ma=mg\sin{\theta}-\mu mg\cos{\theta}=0\\\tan{\theta}=\mu=\frac{1}{3}\\\theta=18.43\degree"